OpenStudy (anonymous):

Can someone pls xplain how they came up with 63 in this equation: 35C=__F F=9/5C +32 F=9/5(35)+32 F=63+32 F=95 So, 35C = 95F thnks for any help

OpenStudy (foolaroundmath):

\[\frac{9}{5}\times 35 = \frac{9}{\cancel{5}1}\times \cancel{35}7 = 9 \times 7 = 63\]

OpenStudy (anonymous):

oh thanks